In an area of the Midwest, records were kept on the relationship between the

rainfall (in inches) and the yield of wheat (bushels per acre).

     Rain (inches)

10.5

8.8

13.4

12.5

18.8

10.3

7.0

15.6

16.0

Yield (bushels/acre)

50.5

46.2

58.8

59.0

82.4

49.2

31.9

76.0

78.8

Which is the residual for y given x = 16.0? Assume that the variables x and y have a significant correlation.
Answer #1

The necessary calculation done below

Observations Rain (X) 1 2 3 4 5 6 7 8 9 sum mean Yield (v) 10.5 8.8 13.4 12.5 18.8 10.3 7 15.6 (x-xbar) 2 (x-xbar)*(y-yba 50.

xbar= mean(x) , ybar=mean(y)

The least square of regression line is given by,

b1= sum((x-xbar)*(y-ybar)) / sum((x-xbar)^2) = 504.13/115.1223 = 4.379085

b0=ybar-b1*xbar = 59.2 - 4.379085 * 12.5444 = 4.266811

The regression equation for predicting yeild is given by,

yhat = b0 + b1*x = 4.266811 + 4.379085* x

predicting yeild for x=16 is given by,

yhat = 4.266811 + 4.379085* x

       = 4.266811 + 4.379085* 16

       = 74.332172

The residual for x=16 is given by,

Residual = Actual value of Y at x=16 - Predicted value of Y at x=16

              = 78.8 - 74.332172

              = 4.5

The residual at x=16 is 4.5